OpenStudy (anonymous):

7c^2=91? How do you solve this?

OpenStudy (anonymous):

7c^2=91 c^2=91/7 c^2=13 \[c=\pm \sqrt{13}\]

OpenStudy (anonymous):

THANK YOU! Now what about n^2=83

OpenStudy (anonymous):

well when I had c^2=13 i took the plus or minus squareroot of 13. what do you think the answer will be for n^2=83?

OpenStudy (anonymous):

n plus or minus the sq rt of 83??

OpenStudy (anonymous):

\[n \pm \sqrt{83}\]

OpenStudy (anonymous):

perfect :D

OpenStudy (anonymous):

what about 5x^2+12=87

OpenStudy (phi):

I'm sure you meant \[ n = \pm \sqrt{83} \]

OpenStudy (anonymous):

ya i did but how do u solve the other one

OpenStudy (phi):

to solve, subtract 12 from both sides. then divide by 5 (both sides) then take the square root (of both sides) what do you get?

OpenStudy (phi):

If this does not make sense, this and the following 2 videos might help http://www.khanacademy.org/video/simple-equations?topic=core-algebra it doesn't do square roots, but you have to walk before you can run.

OpenStudy (anonymous):

[x=\sqrt{15}\] is that right

OpenStudy (phi):

let's check 5x^2+12=87, with x= sqrt(15) 5*sqrt(15)*sqrt(15)+12= 5*15+12= 75+12= 87 yes, it works.

OpenStudy (phi):

all these types of problems follow the same pattern, so try another and see how it goes.